Assume that there is no storage or other carrying costs/value for corn. Assume that corn is not perishable, that is it can be stored and used later and that the continuous compounded risk-free rate is 20% per year.

If a 6-month corn future currently costs $4.75 a bushel and the spot price currently costs $4.50.

Suppose that investors had to pay a short sale fee whenever they short a contract. Assume that this fee was calculated like a carrying cost.

So that if you shorted corn you would have to pay short-selling costs when you enter the shorting contract.

Find the value of short selling costs that is consistent with no arbitrage opportunities.
Answer #1

The Spot price after 6 months = 4.50 + 10% of 4.50

= 4.50+0.45

= 4.95

Short selling cost = 4.75 - 4.95 = (20)